1. y²dx-(xy+y²)dy=0


Let substitution v=y/x ; y=vx ; dy= vdx+xdv



2. x²dy+(2xy-y²)dx=0


Let substitution v=y/x ; y=vx ; dy= vdx+xdv

1
Expert's answer
2022-02-21T05:54:22-0500

1. The given equation can be written as "\dfrac{dy}{dx} = \dfrac{y^2}{xy + y^2}"

Put "y=vx, ~\text{then~} \dfrac{dy}{dx}= v + x\dfrac{dv}{dx}" . The given equation becomes


"\begin{aligned}
v + x\dfrac{dv}{dx} &= \dfrac{v^{2}x^2}{vx^2+v^2x^2}\\
&=\dfrac{v^2}{v+v^2}\\
x\dfrac{dv}{dx} &= \dfrac{v}{1+v} -1\\
&=-\dfrac{v^2}{1+v}\\
\dfrac{1+v}{v^2}dv & = -\dfrac{dx}{x}\\
\dfrac{1}{v^2}dv + \dfrac{1}{v}dv & = -\dfrac{dx}{x}\\
\end{aligned}\\

\text{Integrating both sides, we get}\\
\begin{aligned}
-\dfrac{1}{v} + \log v &= -\log x + \log c\\
-\dfrac{1}{v} &= - \log v -\log x + \log c\\
-\dfrac{1}{v} &= - (\log v + \log x - \log c)\\
-\dfrac{1}{v} &= - \log\left(\frac{vx}{c}\right)\\
\dfrac{x}{y} &= \log\left(\frac{y}{c}\right)~~~(\text{Using $y = vx, v = \frac{y}{x}$})\\
e^{\frac{x}{y}} &= \frac{y}{c}\\
\therefore y &= ce^{\frac{x}{y}}
\end{aligned}"


2. The equation can be written as "\dfrac{dy}{dx} = \dfrac{y^2 - 2xy}{x^2}"

Put "y=vx, ~\text{then~} \dfrac{dy}{dx}= v + x\dfrac{dv}{dx}" . The given equation becomes


"\begin{aligned}
v + x\dfrac{dv}{dx} &= \dfrac{v^{2}x^2-2vx^2}{x^2}\\
&=v^2 - 2v\\
x\dfrac{dv}{dx} &= v^2 - 3v\\
\dfrac{dv}{v(v-3)} & = \dfrac{dx}{x}\\
\end{aligned}\\

\text{Using partial fractions for the term on LHS, we get}\\
\dfrac{1}{3}\left[\dfrac{dv}{v-3} - \dfrac{dv}{v}\right] = \dfrac{dx}{x}\\
\dfrac{dv}{v-3} - \dfrac{dv}{v} = 3\dfrac{dx}{x}\\

\text{Integrating both sides, we get}\\
\begin{aligned}
\log (v-3) - \log v &= 3\log x + \log c\\
\log\left(\dfrac{v-3}{v}\right)&=\log x^3 + \log c\\ 
\log\left(\dfrac{v-3}{v}\right)&=\log (cx^3)\\ 
\dfrac{v-3}{v} &=cx^3~~~(\text{taking anti-logarithms})\\
1 - \frac{3}{v} &=cx^3\\
1 - \frac{3x}{y} &=cx^3~~~(\text{Using $y = vx$})\\
\frac{3x}{y} &= 1 - cx^3\\ 
\therefore y &= \frac{3x}{1 - cx^3}
\end{aligned}"